OpenStudy (anonymous):

I need help with this log function... 2^2x +^x-12=0

OpenStudy (anonymous):

[2^2x +2^x-12=0\]

hartnn (hartnn):

put 2^x = y i think you mean \[2^{2x} +2^x-12=0\]

hartnn (hartnn):

\[(2^{x})^2 +2^x-12=0 \\ 2^x=y \implies y^2+y-12=0\] can you solve this quadratic in y ?

OpenStudy (anonymous):

Didn't think of doing it that way.. I'll try it now

hartnn (hartnn):

ask if you get stuck anywhere..

OpenStudy (anonymous):

I get -7/2 and 3...

hartnn (hartnn):

y^2+y-12=0 (y+4)(y-3)=0 y=-4,3 what you got -7/2 and 3 for ? x ?

OpenStudy (anonymous):

Yup

hartnn (hartnn):

did you get how y=-4,3 ?

OpenStudy (anonymous):

The answer is supposed to be ln3/ln2

OpenStudy (anonymous):

Yes, that part i did.. and i plugged those Y's into X

hartnn (hartnn):

\(2^x=3 \\ \ln 2^x=\ln 3 \\ x \ln 2 = \ln 3 \\ x= \ln3/ \ln 2\)

OpenStudy (anonymous):

Just saw where I messed up.. I pugged the Y's into 2^x... Thanks!!

hartnn (hartnn):

welcome ^_^